# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

### v2

@@ Abstract
Achieving high model robustness under adversarial settings is widely recognized as demanding considerable training samples. Recent works propose semi-supervised adversarial training (SSAT) methods with external unlabeled or synthetically generated data, which are the current state-of-the-art. However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage. In this paper, we propose unlabeled data reduction strategies to improve the efficiency of SSAT. Specifically, we design novel latent clustering-based techniques to select or generate a small critical subset of data samples near the model’s decision boundary. While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting. Comprehensive experiments on benchmark datasets demonstrate that our methods can significantly reduce SSAT’s data requirement and computation costs while preserving its strong robustness advantages. In particular, our latent-space selection scheme based on k-means clustering and our guided DDPM fine-tuning approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data and approximately $4\times$ less total runtime.
@@ I Introduction
Deep neural networks (DNNs) are highly vulnerable to adversarially perturbed inputs [1, 2]. This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness [2, 3, 4, 5]. Among them, adversarial-training-based algorithms are most popular [7, 8, 9]. As illustrated by Schmidt et al. [10], adversarially robust learning is inherently challenging, requiring a much larger sample complexity than standard learning. To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [11, 12] or generated synthetic data [13, 14, 15]. Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store the extra data and a much longer training time for algorithms like adversarial training to converge (see Appendix A for detailed discussions on related work).
@@
Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy. The ultimate goal is to maximize model robustness while utilizing as few additional data samples as possible. Inspired by Zhang et al. [16], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary. Specifically, we introduce three data selection strategies, including PCS, a simple approach that prioritizes unlabeled data points with low prediction confidence, and two advanced latent clustering-based selection schemes: LCS-KM and LCS-GMM, where we identify the most uncertain data in the model’s latent embedding space based on the corresponding clustering technique. The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset. In addition, we propose novel DDPM fine-tuning approaches with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models. Our data reduction schemes balance the ratio of boundary critical points and the remaining non-boundary points to avoid potential overfitting to the boundary distribution. By applying the aforementioned data reduction schemes, the high unlabeled data and computation requirements of SSAT can be significantly alleviated while largely preserving its robustness performance.
@@
Contributions. We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section II). To realize this goal, we propose to optimize the performance of SSAT by refining the model’s decision boundary in the regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points (Section III-A). Specifically, we propose various data selection schemes to identify boundary-adjacent points (Section III-B), including a simple method based on prediction confidence and two advanced selection techniques by clustering the unlabeled or generated data in the latent embedding space. Additionally, we introduce three guided DDPM fine-tuning approaches (Section III-C), aiming to generate the subset of desirable data directly, thus avoiding the computation overhead of pre-generating the whole extra dataset.
@@
Comprehensive experiments on image benchmarks demonstrate that our data selection schemes, especially for the best-performing LCS-KM approach, significantly improve the data and computational efficiency of baseline SSAT methods, while achieving comparable robust accuracies (Section IV-A). For instance, utilizing $10\%$ extra unlabeled data identified by LCS-KM in SSAT can achieve similar robustness performance on SVHN compared to using the entire dataset, while reducing the total runtime to approximately $25\%$. By directly generating the exact amount of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can reduce the computation overhead while maintaining robustness (Section IV-C). Moreover, we demonstrate the generalizability of our data reduction schemes to a real-world COVID-related medical application (Section V-A) and visualize the distributions of selected data points to further illustrate why our LCS-KM method always performs the best (Section V-B). Finally, we provide general guidelines on how to choose the key hyperparameters for maximizing the SSAT performance (Section V-C), and conclude with more discussions on related literature, limitations of our study, and future research directions (Section VI). Our work reveals the importance of prioritizing boundary-adjacent points for improving SSAT’s efficiency and highlights the superiority of latent-based k-means clustering, potentially assisting the development of more efficient and scalable robust learning algorithms for real-world applications.
II Inefficiency of Existing SSAT Methods
II-A Semi-Supervised Adversarial Training
@@
where $\gamma\geq 0$ is a hyperparameter that controls the trade-off between labeled and unlabeled data. A typical approach used in prior work is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\gamma$) but fix the batch size to optimize for the best SSAT performance [13]. For simplicity, we term $\gamma$ as the per-batch ratio. When $\gamma$ is set as $1$, Equation II-A reduces to the optimization objective of vanilla adversarial training (AT) [7]. In Equation II-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss. Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
@@
II-B Limitations of Prior Work
@@
Data Inefficiency. Although incorporating unlabeled data alleviates the sample complexity barriers of adversarially robust learning [10] and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. Such a trend is clearly documented in the leaderboard of RobustBench [17]. For CIFAR-10 image classification tasks, state-of-the-art SSAT algorithms either select 500500K external Tiny Imagenet samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original 5050K CIFAR-10 training images. Specifically, the method proposed by Gowal et al. [13] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [7] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data. Training SSAT algorithms on the huge unlabeled dataset demands significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and carbon footprint. In addition, handling a larger dataset in ML often means increased GPU processing load and memory usage.
@@
High Computation Cost. Apart from the data inefficiency issue, SSAT also requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of total training epochs compared with vanilla AT (see Figure 1 for supporting evidence). Intuitively, learning meaningful and robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more difficult. As our experiments illustrate, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples. Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster. We note that vanilla AT has already been criticized for its high computation [18, 19], due to the iterative PGD steps involved in dealing with the inner maximization problem (Equation 3). SSAT methods incur even higher computations, since they converge more slowly and usually adopt larger model architectures (see Table I for concrete runtime comparisons). Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
II-C Problem Formulation
@@
Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for achieving high model robustness. To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
@@
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation II-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$ (5)
where $\ell_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in Section III-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
@@
where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model. As will be detailed in Section III-C, we adopt a pretrained DDPM [23] as the backbone while strategically fine-tuning it with novel training losses for guided data generations. Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
III Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement. In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section III-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections III-B and III-C).
III-A Prioritize Boundary-Adjacent Unlabeled Data
Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity. Inspired by prior literature [16, 24] that emphasizes the unbalanced data importance for vanilla adversarial training (see Appendix A for more discussions), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose to identify or generate a small set of vulnerable yet valuable unlabeled data points, which are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield more robustness enhancement to adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy function for locating difficult-to-classify data points. Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points. By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.

### v3

@@ Abstract
Learning robust models under adversarial settings is widely recognized as requiring a considerably large number of training samples. Recent work proposes semi-supervised adversarial training (SSAT), which utilizes external unlabeled or synthetically generated data and is currently the state of the art. However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage. In this paper, we propose data reduction strategies to improve the efficiency of SSAT by optimizing the amount of additional data incorporated. Specifically, we design novel latent clustering-based techniques to select or generate a small, critical subset of data samples near the model’s decision boundary. While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points, thereby avoiding overfitting. Comprehensive experiments across image benchmarks demonstrate that our methods can effectively reduce SSAT’s data requirements and computational costs while preserving its strong robustness advantages. In particular, our latent-space selection scheme based on k-means clustering and our guided diffusion-based approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data. When compared to full SSAT trained to convergence, our methods reduce total runtime by approximately $3\times$ to $4\times$ due to strategic prioritization of unlabeled data.
@@ I Introduction
Deep neural networks (DNNs) are shown to be extremely vulnerable to adversarially perturbed inputs [40, 15]. Such a phenomenon has raised serious concerns about the reliability of deploying DNNs in security- or safety-critical applications and has driven numerous research into designing methods to enhance model robustness [15, 31, 3, 2]. Among them, adversarial training and its variants are most popular [27, 43, 50]. As illustrated by Schmidt et al. [34], adversarially robust learning is inherently challenging, requiring a much larger sample complexity than standard learning. To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [5, 49] or generated synthetic data [17, 35, 44]. Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically require vast amounts of additional data, suggesting the need for much larger hardware to store the extra data and a longer training time for algorithms like adversarial training to converge.
@@
Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy. The ultimate goal is to maximize model robustness while utilizing as few additional data samples as possible. Inspired by Zhang et al. [51], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary. Specifically, we introduce three selection strategies, including PCS, a simple approach that prioritizes data points with low prediction confidence, and two advanced latent clustering-based schemes: LCS-KM and LCS-GMM, where we identify the most uncertain region in the model’s latent embedding space using clustering techniques. The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset. In addition, we propose a novel generative method by fine-tuning a denoising diffusion probabilistic model (DDPM) with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models. All our techniques are designed to be lightweight while balancing the ratio of boundary-critical points and the remaining non-boundary points, thereby reducing overall computational costs and avoiding potential overfitting to the boundary distribution. By applying our data reduction techniques, the high requirements for unlabeled data in SSAT can be largely alleviated.
@@
Contributions. We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT algorithms (Section III). To achieve this goal, we propose optimizing SSAT performance by refining the model’s decision boundary in regions of high uncertainty and strategically prioritizing boundary-adjacent data points (Section IV-A). Specifically, we introduce different selection schemes to identify boundary-adjacent points (Section IV-B), including a straightforward method based on prediction confidence and two advanced selection techniques that utilize clustering of unlabeled or generated data in the model’s latent embedding space. In addition, we propose three guided DDPM fine-tuning approaches (Section IV-C) to directly generate a subset of desirable data, thereby avoiding the computational overhead of pre-generating the entire dataset.
@@
Comprehensive experiments on image benchmarks demonstrate that our data selection strategies, particularly for the best-performing LCS-KM approach, significantly improve data efficiency compared to baseline SSAT methods, while achieving comparable test robust accuracies and reducing overall computation costs (Section V-A). For instance, selecting $10\%$ extra DDPM-generated data based on LCS-KM in the SSAT algorithm can achieve robustness performance on SVHN similar to that of using the entire unlabeled dataset. Compared to full SSAT trained to convergence, this reduces the total runtime to approximately $25\%$ due to faster convergence enabled by the strategically selected data. By directly generating the exact number of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can further reduce computational overhead while maintaining robustness (Section V-C). Moreover, we demonstrate the generalizability of our data reduction schemes to a COVID-related medical application task (Section VI-A) and visualize the distributions of selected data points to illustrate why our LCS-KM method consistently performs the best (Section VI-B). We conclude with the limitations of our study and future directions (Section VII). Our work highlights the importance of prioritizing boundary-adjacent points and latent-based k-means clustering in SSAT, potentially assisting in the development of more efficient, robust learning algorithms for real-world applications.
II Related Work
II-A Adversarially Robust Learning
Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [40]. Earlier attacks, such as fast gradient sign method (FGSM) [15] and projected gradient descent (PGD) [27], proposed gradient-based constrained optimization to find adversarial examples. Carlini and Wagner introduced CW attacks to search for minimal but unconstrained perturbations that cause misclassifications [4]. Recent Auto-PGD attacks [10] automate gradient steps to more efficiently produce worst-case perturbations, while AutoAttack [8] is an ensemble of diverse, parameter-free attacks, making them most reliable for evaluating the robustness of DNNs.
Adversarial Training. To attain robustness, adversarial training (AT) and its variants are most popular. For instance, Madry et al. proposed to train DNNs on adversarial inputs using PGD attacks [27], while Zhang et al. introduced TRADES [50], which balances the trade-off between robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting to adversarially perturbed samples produced during training and for their high computational costs, which hinder their deployment in real-world applications. In particular, Rice et al. empirically demonstrated that robust overfitting is a prevalent phenomenon in adversarial training [33], where early stopping is suggested as an effective approach for mitigating this issue. Techniques, such as Free AT [37] and Fast AT [45], offer promising solutions to reduce the computational burden of adversarial training. However, they inevitably suffer from decreased robustness performance.
Moreover, a line of research has pointed out that treating all data points equally in adversarial training is not optimal, as different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [51]. Consequently, they should be given less importance during training to prevent the model from over-optimization, which could negatively affect its generalization on unseen test data. In a similar vein, Hua et al. advocated for a more targeted approach [22], recommending PGD training applied primarily to examples situated near the model’s decision boundaries. The above techniques enhance the model’s adversarial robustness, as perturbations on boundary-adjacent data are more informative in guiding the optimization process of adversarial training. Our work builds on these ideas by strategically prioritizing boundary points from the unlabeled dataset to address the unique challenges associated with SSAT algorithms.
SSAT. Utilizing additional unlabeled data to improve model performance is an active field of research [6]. In the adversarial context, Schmidt et al. [34] highlighted a key observation: robust generalization often necessitates a significantly larger training set size to attain a similar level of performance compared to standard learning. To address the sample complexity challenge, a line of work has proposed semi-supervised adversarial training (SSAT), which utilizes extra unlabeled data, often acquired from a similar but slightly different distribution [5, 1, 29, 49]. In particular, Carmon et al. introduced the robust self-training framework [5], which first trains an intermediate model using available labeled data and then uses the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled dataset, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework that incorporates both labeled and unlabeled data [29], providing formal guarantees on adversarial robustness.
Leveraging synthetic data produced by state-of-the-art generative models to build robust models has been explored in more recent adversarial ML literature [17, 35, 44]. For instance, Gowal et al. employed various unconditional generative models, including DDPM [21], to produce a large synthetic dataset for incorporation in adversarial training [17]. While enhancing robustness, all the aforementioned SSAT methods require a considerable amount of additional unlabeled data. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data, aiming to achieve both efficiency and effectiveness in semi-supervised adversarial training.
II-B Data-Efficient Deep Learning
Dataset Distillation. Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions. For instance, Wang et al. proposed a foundational approach to dataset distillation [42], condensing large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching [52], where distilled data points are optimized to match the gradients of the original dataset. This technique enhances distillation performance, allowing models to learn more effectively from reduced data. In addition, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [39], facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. Coreset selection [36, 24] aims to identify a small, representative subset of data that retains the model performance when trained on it, often relying on diversity-based metrics or optimization techniques to ensure that the selected subset is sufficiently informative. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts [23], while Xia et al. introduced the concept of moderate coreset [47], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [28].
Moreover, Dolatabadi et al. adapted the coreset selection technique to adversarial contexts, offering a task-specific solution to enhance the efficiency of adversarial training [12]. In particular, they proposed adversarial coreset selection, which identifies a compact data subset by minimizing the gradient approximation error, and then utilizes it for adversarial training, with selection taking place at regular intervals. Compared with vanilla AT, the growing reliance on large volumes of generated or external unlabeled data for SSAT presents unique challenges. Due to the significantly expanded training dataset of SSAT, iteratively refining data selection during the entire training process is computationally infeasible. Our work complements the aforementioned methods for efficient SSAT, focusing on developing improved data reduction schemes that do not compromise robustness or scalability.
III Inefficiency of Existing SSAT Methods
III-A Semi-Supervised Adversarial Training
@@
where $\gamma\geq 0$ is a hyperparameter that controls the trade-off between labeled and unlabeled data. A typical approach used in prior work is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\gamma$) but fix the batch size to optimize for the best SSAT performance [17]. For simplicity, we term $\gamma$ as the per-batch ratio. When $\gamma$ is set as $1$, Equation III-A reduces to the optimization objective of vanilla adversarial training (AT) [27]. In Equation III-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss. Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
@@
III-B Limitation of Prior Work
@@
Data Inefficiency. Although incorporating unlabeled data alleviates the sample complexity barriers of adversarially robust learning [34] and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. Such a trend is clearly documented in the leaderboard of RobustBench [8]. For CIFAR-10 image classification tasks, state-of-the-art SSAT algorithms either select 500500K external Tiny ImageNet samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original 5050K CIFAR-10 training images. Specifically, the method proposed by Gowal et al. [17] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [27] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data. Training SSAT algorithms on the vast, unlabeled dataset requires significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and a larger carbon footprint. In addition, handling a larger dataset in machine learning often means increased GPU processing load and memory usage.
@@
High Computation Cost. Apart from the data inefficiency, SSAT requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times training epochs compared with vanilla AT (see Figure 1 for supporting evidence). Intuitively, learning meaningful, robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more challenging. As will be illustrated in our experiments, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples. Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster. We note that vanilla AT has already been criticized for its high computation costs [37, 45], due to the iterative PGD steps involved in solving the inner maximization problem (Equation 3). SSAT algorithms incur even higher computations, as they converge more slowly and often employ larger model architectures (see Table I for concrete runtime comparisons). Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
III-C Problem Formulation
@@
Recognizing the data and computational inefficiencies of SSAT, we propose investigating whether the large size of unlabeled data is necessary for achieving high model robustness. To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
@@
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation III-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right],$ (5)
where $\ell_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in Section IV-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
@@
where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model. As will be detailed in Section IV-C, we adopt a pretrained DDPM [21] as the backbone while strategically fine-tuning it with novel training losses for guided data generations. Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
IV Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement. In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section IV-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections IV-B and IV-C).
IV-A Prioritize Boundary-Adjacent Unlabeled Data
Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity. Inspired by prior literature [51, 22] that emphasizes the unbalanced data importance for adversarial training, we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose identifying or generating a small set of vulnerable yet valuable unlabeled data points that are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield enhanced robustness against adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to provide a meaningful proxy for locating difficult-to-classify data points. Since $f_{\hat{\theta}}$ is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide useful information for identifying these critical points. By focusing more on these critical boundary points, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.